# Quantum Circuit Switching with One-Way Repeaters in Star Networks

### v1

@@ II-B Requests and performance measures
  • The service time distribution is hard to compute analytically. The reason is that calculating the number of packets that need to be sent until $n$ are successfully received over a time window $w$ is a challenging task [19]. When packets are sent in batches of $m$ and each packet can be lost with probability $p$, the number of batches until success is denoted by the random variable $B_{n,w,p,m}$. The mean service time can be computed in terms of the mean value of $B_{n,w,p,m}$:
    $\mathbb{E}\left[T_{\mathrm{service}}\right]=\frac{2L}{c}+t_{\mathrm{fwd}}(2N+1)\mathbb{E}\left[B_{n,w,p,m}\right],$ (3)
    where $c$ is the speed of light in the communication channels. The first term corresponds to the time required for a packet (or a batch of packets) to travel from one user to another. The second term corresponds to the delay introduced by the forwarding stations: each station takes time $t_{\mathrm{fwd}}$ to forward a packet, each packet is forwarded at $12N+$ 1 intermediate repeaters, and a total of $\mathbb{E}\left[B_{n,w,p,m}\right]$ batches of packets must be sent on average to meet the request. In Appendix A, we explain how to compute the probability distribution of $B_{n,w,p,m}$ (in some cases, numerical approximations are required). Note that $m\leq k$, since repeaters can only forward at most $k$ packets simultaneously.
@@ III-A Critical number of users
$\mathbb{E}\left[T_{\mathrm{service}}\right]_{p=1}=\frac{2L}{c}+t_{\mathrm{fwd}}\left\lceil\frac{n}{m}\right\rceil.$ (7)
This expression can be derived using (3) and the fact that $\mathbb{E}[B_{n,w,1,m}]=\left\lceil n/m\right\rceil$ (for further details on the latter, see Appendix A). From (6) and (7), one can show that, for a small budget, the critical number of users is larger when packets are distributed sequentially rather than in parallel. This can be seen in Figure 3, which shows $u_{\mathrm{crit}}$ for increasing number of forwarding stations. It is possible to provide service to a larger number of users (i.e., increase $u_{\mathrm{crit}}$) by increasing the number of forwarding stations $k$, but only when packets are distributed sequentially. In particular, the critical number of users scales as $\sqrt{k}$. When packets are distributed in parallel, increasing the number of stations beyond $k=n$ does not allow for an increased number of users. This is a consequence of the parallel distribution model: we assume that all the forwarding stations are used simultaneously for a single request. When packet distribution is deterministic ($p=1$), only $n$ forwarding stations are needed to distribute $n$ packets in parallel, and increasing $k$ beyond this number will not provide any added benefit – those extra forwarding stations will remain unused.
@@
When the financial budget is large (use case 2), the service time follows a nontrivial distribution and the mean value cannot be computed with (7) anymore. Depending on the values of $p$ and $w$, we were able to compute the mean service time analytically or required Monte Carlo sampling (see Appendix A for further details). Nevertheless, the scaling of $u_{\mathrm{crit}}$ with $k$ remains similar to the behavior shown in Fig. 3 for the small-budget case (we provide the same plot for a large budget and different combinations of $N,L,w$ in Appendix E). We observed that, for most parameter regimes, sequential distribution can still provide service to a larger number of users than parallel distribution. However, this is not true in the parameter regime where resources are scarce, namely, when: ($i$) packet distribution is not deterministic ($p<1$; specifically in our use case, $p\approx 0.7$), ($ii$) applications require many states to be distributed over a short period of time or users have short-lived quantum memories (small $w$ compared to $n$), and ($iii$) there are only a few forwarding stations per repeater (small $k$). In a situation that meets these three conditions, many states must be successfully delivered over a short time window. Since each packet has a nonzero probability of failure, this process can take a very long time when packets are delivered one by one, as successful packets older than the time window are discarded. Hence, parallel distribution can complete requests much faster and therefore provide service to a larger number of users. This reasoning only holds if $k$ is small: if there is a large number of forwarding stations, more users can be served simultaneously with sequential distribution.
@@
Figure 4: Parallel distribution of packets is generally faster. Relative difference in mean sojourn time (MST) between sequential and parallel packet distribution, for different numbers of users $u$ and forwarding stations $k$. (a) Small-budget use case ($p=1$; $w\geq n$; $N=0$, $L=1$ km) and (b) large-budget use case ($p\approx 0.7$) with $N=0$, $L=7.5$ km, and $w=8$. Sequential/parallel distribution provides lower MST in blue/red regions. In regions with an ‘s’/‘p’, only sequential/parallel distribution can provide service (i.e., yield finite MST). In dark regions with an ‘x’, no service is possible. Parameters used in this figure: $n=7$, $\lambda_{0}=10^{-4}$ $\mu$s-1, $c=0.2$ km/$\mu$s, $t_{\mathrm{fwd}}=100$ $\mu$s. MSTs in (a) calculated with (7). MSTs in (b) calculated with Monte Carlo sampling with $10^{7}$ samples (the standard error in the relative difference in MST was below 0.5 for every combination of parameters).
@@
In the large-budget scenario, the asymptotic behavior in $k$ is similar to the small-budget case. The mean service time when $p<1$ is lower bounded by (7), and therefore the critical number of users for $p=1$ (small budget) is an upper bound for $p<1$ (large budget). In particular, this means that $u_{\mathrm{crit}}$ cannot scale faster than $\sqrt{k}$ with sequential distribution, and it converges to a constant as $k\rightarrow\infty$ with parallel distribution. As mentioned earlier, we provide some graphical examples in Appendix E.
@@ III-B Mean sojourn time
  • For a fixed number of forwarding stations $k$, parallel distribution is faster than sequential distribution when the number of users is small. As we increase the number of users (for fixed $k$), the advantage of parallel over sequential distribution generally decreases. For some values of $k$, sequential distribution eventually becomes slightly better. This happens because the main feature of sequential distribution is that multiple requests can be processed simultaneously, and this is particularly beneficial when there is a large number of requests (in our model, a large number of users implies a large number of requests). In some other cases, parallel distribution is always better (e.g., in the small-budget case from Figure 4a with $k>6$). When parallel distribution has a larger critical number of users (e.g., in the large-budget case from Fig 4b with $k=2,3$), the advantage in MST over sequential distribution can actually increase dramatically as the number of users increases (e.g., increasing $u$ from 4 to 5 when $k=2$, in Figure 4b). The reason is that the MST of sequential distribution diverges as we get close to a region where only parallel distribution is possible. This can be seen more clearly in Appendix F, where we plot the MST vs the number of users for fixed $k$.
  • The difference in MST converges to a constant value as $k\rightarrow\infty$. In fact, the MST should converge to a constant as $k\rightarrow\infty$ for both sequential and parallel distribution. Intuitively, once we have enough forwarding stations to meet all incoming requests, there is no benefit in increasing the number of stations per repeater. In Appendix A, we discuss this more formally.
@@ III-C Many users over long distances
Specifically, we want to analyze the critical distance $L_{\mathrm{crit}}$ that overloads the system, i.e., the value of $L$ that makes the load from (4) equal to 1. Combining (3) and (4), we find that the critical distance is given by
$L_{\mathrm{crit}}=\frac{ck}{\lambda_{0}u(u-1)m}-\frac{ct_{\mathrm{fwd}}}{2}(2N+1)\mathbb{E}\left[B_{\mathrm{n,w,p(L_{\mathrm{crit}},N),m}}\right].$ (8)
@@
The scaling with the number of users is therefore upper bounded by $\sim u^{-2}$. This unveils a tradeoff between the number of users and the distances between them. When service is provided to a large number of users, the number of incoming requests will increase and this will put a higher load on the system. If the distances between users are large, service times will increase, since packets will need to travel further away, which will also increase the load. Consequently, we must decrease $u$ to increase $L$ and vice versa. Figure 5 shows an example where we can observe this effect. The plot assumes sequential distribution of packets and specific values of $w$ and $k$, but we have observed the same behavior for parallel distribution and other parameter values, as shown in Appendix G.
We have shown that $L$ and $u$ cannot be scaled up simultaneously when $N$ is fixed, but what if we add more intermediate repeaters to boost the probability of successful packet delivery? Would this allow us to place many users further apart? Intuitively, one may think that this is the case, since a larger number of repeaters $N$ provides a larger success probability $p$ (see (5)), which should decrease the service time. However, this intuition is not always correct, since each repeater introduces a delay $t_{\mathrm{fwd}}$ every time it forwards a quantum data packet. These delays accumulate and increase the service time, despite the gain from a higher $p$. If there are too many repeaters, these forwarding delays may overload the system, preventing requests from being met within a finite amount of time.
When the number of users is small, increasing the number of repeaters is actually beneficial: it allows us to increase the distances between users (i.e., larger $N$ provides larger $L_{\mathrm{crit}}$), as can be seen in Figure 5. When the number of users is large, service is only possible if the number of repeaters is small, that is, $L_{\mathrm{crit}}$ drops to zero quicker for larger $N$. As a consequence, for intermediate values of $u$, there is an finite value of $N$ that maximizes the critical distance. For instance, when there are six users in the example from Figure 5, using $N=2$ repeaters allows the users to be further apart than if we used $N=5$ repeaters.
@@ Appendix A Analytical calculation of the mean waiting time and the mean service time
Figure 6: QCS in a star network can be modeled as an M/G/$s$ queue. Illustration of the queue model of a QCS scheme with (a) sequential distribution of packets ($s=k$) and (b) parallel distribution of packets ($s=1$), in a star network with $k$ forwarding stations per repeater. New requests are submitted at rate $\lambda$. After some control time $t_{\mathrm{control}}$, the request is placed in a first-in first-out queue. After some waiting time $T_{\mathrm{wait}}$, the request leaves the queue and begins being processed. Processing the request takes time $T_{\mathrm{service}}$. The service time distribution is given by the distribution of the number of batches of packets that must be sent until the request is met, $B_{n,w,p,m}$ (see Appendix A-B).
@@
$w\rightarrow\infty$: Closed-form - (A.19) and (A.20) Analytic - (LABEL:eq.Pr_B_winf_multi), (A.32), and (A.33)
@@ A-A Waiting time
with $z\equiv 1-\lambda/(s\mu)$. Refs. [33] and [35] provide empirical evidence that the approximation (A.11) works well when the service time distribution has a small coefficient of variation. We observed empirically that, in the use cases studied in this paper, the coefficient of variation of the service time is small (see Appendix B) and therefore we expect this approximation to be accurate. Note also that the approximation assumes the first two moments of the service time distribution are known.
@@ A-B Service time
As explained in the main text, the service time can be computed as
$T_{\mathrm{service}}=\frac{2L}{c}+t_{\mathrm{fwd}}(2N+1)B_{n,w,p,m},$ (A.13)
where $L$ is the distance between each user and the central repeater, $c$ is the speed of light in the physical channels, $t_{\mathrm{fwd}}$ is the forwarding time, and $N$ is the number of repeaters between each user and the central repeater. Recall that $B_{n,w,p,m}$ is the number of batches of quantum data packets that must be sent from user to user until $n$ packets are successfully distributed within a time window $w$, assuming each packet has a success probability $p$ and each batch contains $m$ packets. The mean service time is then given by
$\mathbb{E}\left[T_{\mathrm{service}}\right]=\frac{2L}{c}+t_{\mathrm{fwd}}(2N+1)\mathbb{E}\left[B_{n,w,p,m}\right],$ (A.14)
since $B_{n,w,p,m}$ is the only random variable involved. As shown in Section A-A, the second moment of the service time is also required to compute the waiting time. The second moment can be computed as
$\mathbb{E}\left[T_{\mathrm{service}}^{2}\right]=\left(\frac{2L}{c}\right)^{2}+\frac{4L}{c}t_{\mathrm{fwd}}(2N+1)\mathbb{E}\left[B_{n,w,p,m}\right]+\left(t_{\mathrm{fwd}}(2N+1)\right)^{2}\mathbb{E}\left[B_{n,w,p,m}^{2}\right].$ (A.15)
@@
It is possible to calculate the probability distribution of $B_{n,w,p,m}$ analytically when there is no multiplexing ($m=1$) [19]. When $p=1$, the solution is trivial since we deterministically need $n$ batches to successfully deliver $n$ packets (there is only one per batch). In that case, we have
@@
As shown in ref. [19], in the case of an infinite window ($w\rightarrow\infty$), $B_{n,w,p,1}$ follows a negative binomial distribution, which has the following first two moments:
@@
In the multiplexed case ($m>1$), there is no known analytical solution for every combination of $n$, $w$, $p$, and $m$, to the best of our knowledge. Next, we solve the problem for two cases: ($i$) $p=1$ and ($ii$) $w\rightarrow\infty$. In every other case, we employed Monte Carlo sampling to estimate the probability distribution of $B_{n,w,p,m}$ when needed. First, the case $p=1$ is again trivial. When packets are always successfully delivered, we need $\lceil n/m\rceil$ batches to deliver $n$ packets. Hence,
@@
The survival function can then be used to calculate the first two moments of $B_{n,\infty,p,m}$ as follows:

### v2

@@ II-B Requests and performance measures
  • The service time distribution is hard to compute analytically. The reason is that calculating the number of packets that need to be sent until $n$ are successfully received over a time window $w$ is a challenging task [19]. When packets are sent in batches of $m$ (with $m\leq k$, since repeaters can only forward at most $k$ packets simultaneously) and each packet can be lost with probability $p$, the number of batches until success is denoted by the random variable $B_{n,w,p,m}$. The mean service time can be computed in terms of the mean value of $B_{n,w,p,m}$:
    $\mathbb{E}\left[T_{\mathrm{service}}\right]=\frac{2L}{c}+t_{\mathrm{fwd}}\left(2N+\mathbb{E}\left[B_{n,w,p,m}\right]\right),$ (3)
    where $c$ is the speed of light in the communication channels. The first term corresponds to the time required for a packet (or a batch of packets) to travel from one user to another. The second term accounts for the total delay introduced by the forwarding stations, which depends on the number of repeaters, $N$, and the expected number of batches required, $\mathbb{E}\left[B_{n,w,p,m}\right]$. In Appendix A-B, we provide a derivation of (3), and we also explain how to compute the probability distribution of $B_{n,w,p,m}$ (in some cases, numerical approximations are required).
@@ III-A Critical number of users
$\mathbb{E}\left[T_{\mathrm{service}}\right]_{p=1,N=0}=\frac{2L}{c}+t_{\mathrm{fwd}}\left\lceil\frac{n}{m}\right\rceil.$ (7)
This expression can be derived using (3) and the fact that $\mathbb{E}[B_{n,w,1,m}]=\left\lceil n/m\right\rceil$ (for further details on the latter, see Appendix A-B). From (6) and (7), one can show that, for a small budget, the critical number of users is larger when packets are distributed sequentially rather than in parallel. This can be seen in Figure 3, which shows $u_{\mathrm{crit}}$ for increasing number of forwarding stations. It is possible to provide service to a larger number of users (i.e., increase $u_{\mathrm{crit}}$) by increasing the number of forwarding stations $k$, but only when packets are distributed sequentially. In particular, the critical number of users scales as $\sqrt{k}$. When packets are distributed in parallel, increasing the number of stations beyond $k=n$ does not allow for an increased number of users. This is a consequence of the parallel distribution model: we assume that all the forwarding stations are used simultaneously for a single request. When packet distribution is deterministic ($p=1$), only $n$ forwarding stations are needed to distribute $n$ packets in parallel, and increasing $k$ beyond this number will not provide any added benefit – those extra forwarding stations will remain unused.
@@
When the financial budget is large (use case 2), the service time follows a nontrivial distribution and the mean value cannot be computed with (7) anymore. Depending on the values of $p$ and $w$, we were able to compute the mean service time analytically or required Monte Carlo sampling (see Appendix A for further details). Nevertheless, the scaling of $u_{\mathrm{crit}}$ with $k$ remains similar to the behavior shown in Fig. 3 (we provide the same plot for a large budget and different combinations of $N,L,w$ in Appendix E). We observed that, for most parameter regimes, sequential distribution can still provide service to a larger number of users than parallel distribution. However, this is not true in the parameter regime where resources are scarce, namely, when: ($i$) packet distribution is not deterministic ($p<1$; specifically in our use case, $p\approx 0.7$), ($ii$) applications require many states to be distributed over a short period of time or users have short-lived quantum memories (small $w$ compared to $n$), and ($iii$) there are only a few forwarding stations per repeater (small $k$). In a situation that meets these three conditions, many states must be successfully delivered over a short time window. Since each packet has a nonzero probability of failure, this process can take a very long time when packets are delivered one by one, as successful packets older than the time window are discarded. Hence, parallel distribution can complete requests much faster and therefore provide service to a larger number of users. This reasoning only holds if $k$ is small: if there is a large number of forwarding stations, more users can be served simultaneously with sequential distribution.
@@
Figure 4: Parallel distribution of packets is generally faster. Relative difference in mean sojourn time (MST) between sequential and parallel packet distribution, for different numbers of users $u$ and forwarding stations $k$. (a) Small-budget use case ($p=1$; $w\geq n$; $N=0$, $L=1$ km) and (b) large-budget use case ($p\approx 0.7$) with $N=0$, $L=7.5$ km, and $w=8$. Sequential/parallel distribution provides lower MST in blue/red regions. In regions with an ‘s’/‘p’, only sequential/parallel distribution can provide service (i.e., yield finite MST). In dark regions with an ‘x’, no service is possible. Parameters used in this figure: $n=7$, $\lambda_{0}=10^{-4}$ $\mu$s-1, $c=0.2$ km/$\mu$s, $t_{\mathrm{fwd}}=100$ $\mu$s. MSTs in (a) calculated with (7). MSTs in (b) calculated with Monte Carlo sampling with $10^{4}$ samples (the standard error in the relative difference in MST was below 0.5 for every combination of parameters).
@@
In the large-budget scenario, the asymptotic behavior in $k$ is similar to the small-budget case. When $p<1$, the mean service time is lower bounded by (7), and therefore the critical number of users for $p=1$ (small budget) is an upper bound for $p<1$ (large budget). In particular, this means that $u_{\mathrm{crit}}$ cannot scale faster than $\sqrt{k}$ with sequential distribution, and it converges to a constant as $k\rightarrow\infty$ with parallel distribution. As mentioned earlier, we provide some graphical examples in Appendix E.
@@ III-B Mean sojourn time
  • For a fixed number of forwarding stations $k$, parallel distribution is faster than sequential distribution when the number of users is small. As we increase the number of users (for fixed $k$), the advantage of parallel over sequential distribution generally decreases. For some values of $k$, sequential distribution eventually becomes slightly better. This happens because the main feature of sequential distribution is that multiple requests can be processed simultaneously, and this is particularly beneficial when there is a large number of requests (in our model, a large number of users implies a large number of requests). In some other cases, parallel distribution is always better (e.g., in the small-budget example from Figure 4a with $k>6$). When parallel distribution has a larger critical number of users (e.g., in the large-budget case from Fig 4b with $k=2,3$), the advantage in MST over sequential distribution can actually increase dramatically as the number of users increases (e.g., increasing $u$ from 4 to 5 when $k=2$, in Figure 4b). The reason is that the MST of sequential distribution diverges as we get close to a region where only parallel distribution is possible. This can be seen more clearly in Appendix F, where we plot the MST vs the number of users for fixed $k$.
  • The difference in MST converges to a constant value as $k\rightarrow\infty$. In fact, the MST should converge to a constant as $k\rightarrow\infty$ for both sequential and parallel distribution. Intuitively, once we have enough forwarding stations to meet all incoming requests, there is no benefit in increasing the number of stations per repeater. In Appendix A-C, we discuss this more formally.
@@ III-C Many users over long distances
Specifically, we want to analyze the critical distance $L_{\mathrm{crit}}$ that overloads the system, i.e., the value of $L$ that makes the load from (4) equal to 1. Combining (3) and (4) and enforcing $L\geq 0$, we find that the critical distance is given by
$\begin{split}L_{\mathrm{crit}}=\max\bigg{(}0,\frac{ck}{\lambda_{0}u(u-1)m}-\frac{ct_{\mathrm{fwd}}}{2}\Big{(}2N\\ +\mathbb{E}\left[B_{\mathrm{n,w,p(L_{\mathrm{crit}},N),m}}\right]\Big{)}\bigg{)}.\end{split}$ (8)
@@
The scaling with the number of users is therefore upper bounded by $\sim u^{-2}$. This unveils a tradeoff between the number of users and the distances between them. When service is provided to a large number of users, the number of incoming requests will increase and this will put a higher load on the system. If the distances between users are large, service times will increase, since packets will need to travel further away, which will also increase the load. Consequently, we must decrease $u$ to increase $L$ and vice versa. Figure 5 shows an example where we can observe this effect.
We have shown that $L$ and $u$ cannot be scaled up simultaneously when $N$ is fixed, but what if we add more intermediate repeaters to boost the probability of successful packet delivery? Would this allow us to place many users further apart? Intuitively, one may think that this is the case, since a larger number of repeaters $N$ provides a larger success probability $p$ (see (5)). Increasing $p$ means that we will need to send less (batches of) packets per request, i.e., $\mathbb{E}[B_{n,w,p,m}]$ will be smaller, and consequently the service time (3) should decrease and the critical distance (8) should increase. However, this intuition is not always correct, since each additional repeater introduces an additional delay $t_{\mathrm{fwd}}$, which directly impacts the service time and the critical distance: both (3) and (8) include a linear term in $N$. These linear terms yield an interesting behavior: the delays introduced by the repeaters accumulate and the service time eventually increases (and the critical distance decreases) when increasing the number of repeaters $N$, despite the benefit from a larger $p$. If there are too many repeaters, these forwarding delays may overload the system, preventing requests from being met within a finite amount of time.
When the number of users is small, increasing the number of repeaters is actually beneficial: it allows us to increase the distances between users (i.e., larger $N$ provides larger $L_{\mathrm{crit}}$), as can be seen in Figure 5. When the number of users is large, service is only possible if the number of repeaters is small. That is, when increasing the number of users, $L_{\mathrm{crit}}$ drops to zero quicker for larger $N$. As a consequence, for intermediate values of $u$, there is an finite value of $N$ that maximizes the critical distance. For instance, when there are ten users in the example from Figure 5, using $N=5$ repeaters allows the users to be further apart than if we used $N=10$ repeaters. Note that Figure 5 assumes sequential distribution of packets, although the previous discussion remains general and also applies to parallel distribution, as shown in Appendix G.
@@ IV Outlook
Code availability
The code used to generate all the plots in this paper can be found in the following GitHub repository: https://github.com/AlvaroGI/quantum-circuit-switching.
@@ Appendix A Analytical calculation of the mean waiting time and the mean service time
Figure 6: QCS in a star network can be modeled as an M/G/$s$ queue. Illustration of the queue model of a QCS scheme with (a) sequential distribution of packets ($s=k$) and (b) parallel distribution of packets ($s=1$), in a star network with $k$ forwarding stations per repeater. New requests are submitted at rate $\lambda$. After some control time $t_{\mathrm{control}}$, the request is placed in a first-in first-out queue. After some waiting time $T_{\mathrm{wait}}$, the request leaves the queue and begins being processed. Processing the request takes time $T_{\mathrm{service}}$. Sequential distribution (a) can process up to $k$ requests in parallel, while parallel distribution (b) can only process one at a time. The service time distribution is given by the distribution of the number of batches of packets that must be sent until the request is met, $B_{n,w,p,m}$ (see Appendix A-B).
@@
$w\rightarrow\infty$: Closed-form - (A.19) and (A.20) Analytical - (LABEL:eq.Pr_B_winf_multi), (A.32), and (A.33)
@@ A-A Waiting time
with $z\equiv 1-\lambda/(s\mu)$. Refs. [33] and [35] provide empirical evidence that the approximation (A.11) works well when the service time distribution has a small squared coefficient of variation $C_{\mathrm{service}}^{2}$. We observed empirically that, in the use cases studied in this paper, $C_{\mathrm{service}}^{2}$ is indeed small (see Appendix B) and therefore we expect this approximation to be accurate. Note also that the approximation assumes the first two moments of the service time distribution are known.
@@ A-B Service time
As introduced in the main text, $B_{n,w,p,m}$ is the number of batches of quantum data packets that must be sent from user to user until $n$ packets are successfully distributed within a time window $w$, assuming each packet has a success probability $p$ and each batch contains $m$ packets. The service time can be computed as
$T_{\mathrm{service}}=\frac{2L}{c}+t_{\mathrm{fwd}}(2N+1)+t_{\mathrm{fwd}}\left(B_{n,w,p,m}-1\right),$ (A.13)
where $L$ is the distance between each user and the central repeater, $c$ is the speed of light in the physical channels, $t_{\mathrm{fwd}}$ is the forwarding time, and $N$ is the number of repeaters between each user and the central repeater. The first term corresponds to the time required for the first packet (or batch of packets) to travel from one user to another, without considering forwarding delays. The second term corresponds to the delay introduced by the forwarding stations: each station takes time $t_{\mathrm{fwd}}$ to forward a packet, and the packet (or batch of packets) is forwarded at $12N+$ 1 intermediate repeaters. Therefore, the first two terms account for the total time since the first packet (or batch of packets) is sent from one user until it is received by the other user. The third term accounts for the time it takes to receive the remaining $B_{n,w,p,m}-1$ (batches of) packets: a packet (or batch of packets) is sent every $t_{\mathrm{fwd}}$ units of time, to ensure it will arrive at the next repeater when the previous packet is already processed and leaving the repeater. The mean service time is then given by
$\mathbb{E}\left[T_{\mathrm{service}}\right]=\frac{2L}{c}+t_{\mathrm{fwd}}(2N+\mathbb{E}\left[B_{n,w,p,m}\right]),$ (A.14)
since $B_{n,w,p,m}$ is the only random variable involved. As shown in Section A-A, the second moment of the service time is also required to compute the expected waiting time $\mathbb{E}[T_{\mathrm{wait}}]$. The second moment can be computed as
$\mathbb{E}\left[T_{\mathrm{service}}^{2}\right]=\left(\frac{4L^{2}}{c^{2}}+\frac{8L}{c}t_{\mathrm{fwd}}N+4t_{\mathrm{fwd}}^{2}N^{2}\right)+\left(\frac{4L}{c}t_{\mathrm{fwd}}+4t_{\mathrm{fwd}}^{2}N\right)\mathbb{E}\left[B_{n,w,p,m}\right]+t_{\mathrm{fwd}}^{2}\mathbb{E}\left[B_{n,w,p,m}^{2}\right].$ (A.15)
@@
It is possible to calculate the probability distribution of $B_{n,w,p,m}$ analytically when there is no multiplexing ($m=1$) [19]. When $m=1$ and $p=1$, the solution is trivial since we deterministically need $n$ batches to successfully deliver $n$ packets (there is only one per batch). In that case, we have
@@
As shown in ref. [19], in the case of $m=1$ and an infinite window ($w\rightarrow\infty$), $B_{n,w,p,1}$ follows a negative binomial distribution, which has the following first two moments:
@@
In the multiplexed case ($m>1$), there is no known analytical solution yet for every combination of $n$, $w$, $p$, and $m$, to the best of our knowledge. Next, we solve the problem for two cases: ($i$) $p=1$ and ($ii$) $w\rightarrow\infty$. In every other case, we employed Monte Carlo sampling to estimate the probability distribution of $B_{n,w,p,m}$ when needed. First, the case $p=1$ is again trivial. When packets are always successfully delivered, we need $\lceil n/m\rceil$ batches to deliver $n$ packets. Hence,
@@
The survival function (LABEL:eq.Pr_B_winf_multi) can then be used to calculate the first two moments of $B_{n,\infty,p,m}$ as follows: